zarkam21 (zarkam21):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. 13x-2y=70 2x-2y=4

DivineSolar (divinesolar):

How many more of these do you have?

zarkam21 (zarkam21):

A couple I just don't understand it

DivineSolar (divinesolar):

Oh well look back on what i taught you on your last problem.

DivineSolar (divinesolar):

...No

DivineSolar (divinesolar):

Let us try this again.

zarkam21 (zarkam21):

Is it 11x+70=4

DivineSolar (divinesolar):

no xD

zarkam21 (zarkam21):

DO you set the two equations equal to each other

DivineSolar (divinesolar):

Try solving that.

DivineSolar (divinesolar):

Let me know what you get.

zarkam21 (zarkam21):

13x+2y=70+4y

DivineSolar (divinesolar):

Close, it would be 13x=2y+70, Look it over and let me know if you see how i got that.

zarkam21 (zarkam21):

You got that by cancelling out the 2y right?

DivineSolar (divinesolar):

Yes the 2y's do cancel out :).

zarkam21 (zarkam21):

Yes thats how you got that could you explain how you started did you set the equations equal to each other

DivineSolar (divinesolar):

No i did not were just solving for x right now so the first equation.

DivineSolar (divinesolar):

So now with the given equation -> 13x=2y+70, divide both sides by 13

DivineSolar (divinesolar):

Remember always let me know what you get ^-^.

zarkam21 (zarkam21):

x=2y+10

DivineSolar (divinesolar):

Actually it would be, x= 2/13y + 70/13

DivineSolar (divinesolar):

Now here come's the hard part i will do most of the work, you will do some but here.

DivineSolar (divinesolar):

After that png i want you to, Divide both sides by (-22)/13

zarkam21 (zarkam21):

I got a decimal for that answer

DivineSolar (divinesolar):

Round the decimal to a whole number and let me know what it is.

zarkam21 (zarkam21):

-1.70

zarkam21 (zarkam21):

so 2?

DivineSolar (divinesolar):

Hmm not sure how you got that, and i will see how you are gonna get confused with my answer, however it is actually y = 4

zarkam21 (zarkam21):

Shouldn't it be -4

DivineSolar (divinesolar):

Now you need to do the work in this png :).

DivineSolar (divinesolar):

No ma'am :)

DivineSolar (divinesolar):

Division make's the answer a positive.

zarkam21 (zarkam21):

Ohhh okay =)

zarkam21 (zarkam21):

meaning x=5

DivineSolar (divinesolar):

Very close.

zarkam21 (zarkam21):

soo it would be (5,4)

DivineSolar (divinesolar):

It would actually be x = 6, look it over and please try and see how to get it :).

zarkam21 (zarkam21):

ohhh yes that is correct sooo (6,4)

DivineSolar (divinesolar):

Yes :).

DivineSolar (divinesolar):

x=6 and y=4

DivineSolar (divinesolar):

^-^.

zarkam21 (zarkam21):

Thank you soooooooooo much once again

DivineSolar (divinesolar):

Sure thing.

DivineSolar (divinesolar):

Please close this question :)!